8.) 2.) A bomb calorimeter, or a constant volume calorimeter, is a device often used to determine the heat of combustion of fuels and the energy content of foods.


In an experiment, a 1.0860 g sample of glutaric acid (C5H8O4) is burned completely in a bomb calorimeter. The calorimeter is surrounded by 1.143×103 g of water. During the combustion the temperature increases from 21.61 to 24.66 °C. The heat capacity of water is 4.184 J g-1°C-1.

The heat capacity of the calorimeter was determined in a previous experiment to be 761.1 J/°C.

Assuming that no energy is lost to the surroundings, calculate the molar heat of combustion of glutaric acid based on these data.

C5H8O4(s) + 5O2(g) = 4 H2O(l) + 5 CO2(g) + Energy
Answer #1

there is no heat loss so we can say that all heat released in combustion is absorbed by water and calorimeter.

heat absorbed by calorimeter = heat capacity of calorimeter \times change in temperature of calorimeter

= 761.1 ( J/ oC) \times (24.66 - 21.61)

= 2321.355 J

heat absorbed by water = mass of water \times heat capacity of water \times change in temperature of water

= 1.143 \times 103(g) \times 4.184 ( J K-1 oC-1) \times ( 24.66 - 21.61)

= 14586.05 J

we know that heat released in the combustion = heat absorbed by water + heat absorbed by calorimeter

= 14586.051 + 2321.355

  heat released in the combustion = 16907.406 J

this heat is released by combustion of 1.0860 g of glutaric acid.

convert this weight in mol, molar mass of glutaric acid is 132. so mol = 1.0860/132 = 8.2272 \times 10-3 mol

so heat released by combustion of 8.2272 \times 10-3 mol of glutaric acid. = 16907.406 J

then heat released by 1 mol of glutaric acid = 2055.043 KJ/mol

molar heat of combustion = 2055.043 KJ/mol